How many grams are in 7.4 moles of C6H12O6?


1
Expert's answer
2021-05-10T01:21:25-0400

Molar mass of "C_6H_{12}O_6=180.16g\/mol"


So Number of moles "=\dfrac{\text{ Given mass}}{\text{molar mass}}"


"7.4=\dfrac{x}{180.16}\Rightarrow x=180.16\times 7.4=1333.184 g"


Hence Mass is 1333.184g